OpenStudy (anonymous):

(3xy+3xz)/(4x^2y+4x^2z)

OpenStudy (anonymous):

Are we ask to simplify ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 3/4x, but he counted it wrong

OpenStudy (anonymous):

Factor out the denominator

OpenStudy (anonymous):

4x^2(y+z)

OpenStudy (anonymous):

now separate the fraction\[\frac{ 3xy }{ 4x ^{2}(y+z)}+\frac{ 3xz }{ 4x ^{2}(y+z) }\]

OpenStudy (anonymous):

look at the first part of the problem, what are the like terms? Is there any way to cancel out those like terms?

OpenStudy (anonymous):

the x right?

OpenStudy (anonymous):

thats one of them, there is another

OpenStudy (anonymous):

the y

OpenStudy (anonymous):

yes, now we need to cancel those out in the numerator and denominator. So to do that we need to factor out the y.\[\frac{ 3xy }{ 4x ^{2}y(1-z) }\]

OpenStudy (anonymous):

once you cancel these what would the fraction look like?

OpenStudy (anonymous):

keep in mind to divide exponents you subtract.

OpenStudy (anonymous):

3/4x-z ?

OpenStudy (anonymous):

almost, you would subtract the denominator exponent from the numerator.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

\[x ^{1-2}=x ^{-1}\]

OpenStudy (anonymous):

so 3/4x^-1 -z ? or would it become a -4x?

OpenStudy (anonymous):

\[\frac{ 3 }{ 4x ^{-1}(1-z) }\]\[\frac{ 3 }{ 4x ^{-1}-4x ^{-1}z }\]

OpenStudy (anonymous):

would you factor out the denom?

OpenStudy (anonymous):

We foiled the denom. To factor it would leave you with what we got after canceling the like terms, so we would leave it like that for the moment, and do the same to the second part of the fraction.

OpenStudy (anonymous):

so would it be : 3/4x^-1y?

OpenStudy (anonymous):

it would be \[\frac{ 3 }{ 4x ^{-1}-4x ^{-1}y }\] giving you\[\frac{ 3 }{ 4x ^{-1}-4x ^{-1}z }+\frac{ 3 }{ 4x ^{-1}y }\] What do you think the next step would be? Also a side note in case you didn't already know is that underneath the text box are blue buttons, the equation button lets you enter equations instead of having to use symbols like ^ to represent exponents. The draw button lets you draw, add graphs, tables, ect. then the attach file lets you attach a file. and the post lets others see what you wrote.

OpenStudy (anonymous):

sorry I left out a part of the equation, \[\frac{ 3 }{ 4x^{-1} -4x ^{-1}z}+\frac{ 3 }{ 4x ^{-1}-4x ^{-1} y}\]

OpenStudy (anonymous):

\[\frac{ 3(-4x ^{-1}z) }{4x^{-1}y-4x^{-1}z }\]

OpenStudy (anonymous):

you need to make common denominators by cross multiplying\[\frac{ 3(4x ^{-1}-4x ^{-1}y) }{4x ^{-1}-4x ^{-1}z(4x ^{-1}-4x ^{-1}y) }+\frac{ 3(4x ^{-1}-4x ^{-1}z) }{4x ^{-1}-4x ^{-1}y(4x ^{-1}-4x ^{-1}z) }\]

OpenStudy (anonymous):

do you multiply the the denominators?

OpenStudy (anonymous):

yes, im calculating it right now, so give it a go, and we'll see if we match results

OpenStudy (anonymous):

alright, I got an answer that cannot possibly be the answer. So if you'd like I can look at it, see where I went wrong, and back to you when I can. Sometimes I need to step away for a while and come back with a totally new perspective. So if you have the time, that's what I can do for you. If not, I recommend re-posting the question and having some one else help you out.

OpenStudy (anonymous):

Sorry about that :(

OpenStudy (anonymous):

ok yeah i got lost as well.

OpenStudy (anonymous):

alright, I'll mail you when I got it figured out.

OpenStudy (anonymous):

okay thank you!